# Jumping Ahead: Improving Reconstruction Fidelity with JumpReLU Sparse Autoencoders

### v1

@@ 1 Introduction
In this paper, we introduce JumpReLU SAEs – a small modification of the original, ReLU-based SAE architecture (Ng, 2011) where the SAE encoder’s ReLU activation function is replaced by a JumpReLU activation function (Erichson et al. (2019), previously named TRec by Konda et al. (2015)), which zeroes out pre-activations below a positive threshold (see Fig. 1). Moreover, we train JumpReLU SAEs using a loss function that is simply the weighted sum of a L2 reconstruction error term and a L0 sparsity penalty, eschewing easier-to-train proxies to L0, such as L1, and avoiding the need for auxiliary tasks to train the threshold.
@@ 3 JumpReLU SAEs
In these expressions, $K$ can be any valid kernel function (see Section 2) – i.e. it needs to satisfy the properties of a centered, finite-variance probability density function. In our experiments, we use the rectangle function, $\text{rect}(z):=H\left(z+\tfrac{1}{2}\right)-H\left(z-\tfrac{1}{2}\right)$ as our kernel; however we expect similar results can be obtained with other common kernels, such as the triangular, Gaussian or Epanechnikov kernel. As we show in Section 4, the hyperparameter $\varepsilon$ plays the role of a KDE bandwidth, and needs to be selected accordingly: too low and gradient estimates become too noisy, too high and estimates become too biased. [^5] For the experiments in this paper, we swept this parameter and found $\varepsilon=0.001$ (assuming a dataset normalised such that $\mathbb{E}_{\mathbf{x}}[\mathbf{x}^{2}]=1$) works well across different models, layers and sites. However, we suspect there are more principled ways to determine this parameter, borrowing from the literature on KDE bandwidth selection.
Having defined these pseudo-derivatives, we train JumpReLU SAEs as we would any differentiable model, by computing the gradient of the loss function in Eq. 9 over batches of data (remembering to apply these pseudo-derivatives in the backward pass), and sending the batch-wise mean of these gradients these to the optimiser in order to compute parameter updates.
In Appendix I we provide pseudocode for the JumpReLU SAE’s forward pass, loss function and for implementing the straight-through-estimators defined in Eq. 11 and Eq. 12 in an autograd framework like Jax (Bradbury et al., 2018) or PyTorch (Paszke et al., 2019).
@@ 5 Evaluation
All SAEs used in these evaluations were trained over 8 billion tokens; by this point, they had all converged, as confirmed by inspecting their training curves. See Appendix H for further details of our training methodology.
@@ 7 Discussion
  • JumpReLU SAEs introduce new hyperparameters – namely the initial value of ${\boldsymbol{\theta}}$ and the bandwidth parameter $\varepsilon$ – that require selecting. In practice, we find that, with dataset normalization in place, the default hyperparameters used in our experiments (Appendix H) transfer quite reliably to other models, sites and layers. Nevertheless, there may be more principled ways to choose these hyperparameters, for example by adopting approaches to automatically selecting bandwidths from the literature on kernel density estimation.
@@ 8 Acknowledgements
We thank Lewis Smith for reviewing the paper, including checking its mathematical derivations, and for valuable contributions to the SAE training codebase. We are also grateful to Rohin Shah and Anca Dragan for their sponsorship and support during this project.
@@ Appendix G Additional benchmarking results
Appendix H Further details on our training methodology
@@
Appendix I Pseudo code for implementing and training JumpReLU SAEs
@@
### Implementation of Heaviside step function with custom backward
@@
def heaviside(x):
return (x > 0).astype(x.dtype)
def heaviside_fwd(x):
out = heaviside(x)
cache = x # Saved for use in the backward pass
@@
def heaviside_bwd(cache, output_grad):
x = cache
x_grad = (1.0 / bandwidth) * rectangle(x / bandwidth) * output_grad
return (x_grad,) # A tuple of gradients, one for each input
heaviside.defvjp(heaviside_fwd, heaviside_bwd)
@@
x_grad = zeros_like(x) # We don’t apply STE to x input
@@
l0 = sum(heaviside(feature_magnitudes - threshold), axis=-1)

### v2

@@ 1 Introduction
In this paper, we introduce JumpReLU SAEs – a small modification of the original, ReLU-based SAE architecture (Ng, 2011) where the SAE encoder’s ReLU activation function is replaced by a JumpReLU activation function (Erichson et al., 2019), which zeroes out pre-activations below a positive threshold (see Fig. 1). Moreover, we train JumpReLU SAEs using a loss function that is simply the weighted sum of a L2 reconstruction error term and a L0 sparsity penalty, eschewing easier-to-train proxies to L0, such as L1, and avoiding the need for auxiliary tasks to train the threshold.
@@ 3 JumpReLU SAEs
In these expressions, $K$ can be any valid kernel function (see Section 2) – i.e. it needs to satisfy the properties of a centered, finite-variance probability density function. In our experiments, we use the rectangle function, $\text{rect}(z):=H\left(z+\tfrac{1}{2}\right)-H\left(z-\tfrac{1}{2}\right)$ as our kernel; however similar results can be obtained with other common kernels, such as the triangular, Gaussian or Epanechnikov kernel (see Appendix H). As we show in Section 4, the hyperparameter $\varepsilon$ plays the role of a KDE bandwidth, and needs to be selected accordingly: too low and gradient estimates become too noisy, too high and estimates become too biased. [^5] For the experiments in this paper, we swept this parameter and found $\varepsilon=0.001$ (assuming a dataset normalised such that $\mathbb{E}_{\mathbf{x}}[\mathbf{x}^{2}]=1$) works well across different models, layers and sites. However, we suspect there are more principled ways to determine this parameter, borrowing from the literature on KDE bandwidth selection.
Having defined these pseudo-derivatives, we train JumpReLU SAEs as we would any differentiable model, by computing the gradient of the loss function in Eq. 9 over batches of data (remembering to apply these pseudo-derivatives in the backward pass), and sending the batch-wise mean of these gradients to the optimiser in order to compute parameter updates.
In Appendix J we provide pseudocode for the JumpReLU SAE’s forward pass, loss function and for implementing the straight-through-estimators defined in Eq. 11 and Eq. 12 in an autograd framework like Jax (Bradbury et al., 2018) or PyTorch (Paszke et al., 2019).
@@ 5 Evaluation
All SAEs used in these evaluations were trained over 8 billion tokens; by this point, they had all converged, as confirmed by inspecting their training curves. See Appendix I for further details of our training methodology.
@@ 6 Related work
The activation function defined in Eq. 4 was named JumpReLU in Erichson et al. (2019), although it appears in earlier work, such as the TRec function in Konda et al. (2015). Both TopK and JumpReLU activation functions are closely related to activation pruning techniques such as ASH (Djurisic et al., 2023).
@@ 7 Discussion
  • JumpReLU SAEs introduce new hyperparameters – namely the initial value of ${\boldsymbol{\theta}}$ and the bandwidth parameter $\varepsilon$ – that require selecting. In practice, we find that, with dataset normalization in place, the default hyperparameters used in our experiments (Appendix I) transfer quite reliably to other models, sites and layers. Nevertheless, there may be more principled ways to choose these hyperparameters, for example by adopting approaches to automatically selecting bandwidths from the literature on kernel density estimation.
@@ 8 Acknowledgements
We thank Lewis Smith for reviewing the paper, including checking its mathematical derivations, and for valuable contributions to the SAE training codebase. We also thank Tom Conerly for pointing out errors in an earlier version of Appendix J. Finally, we are grateful to Rohin Shah and Anca Dragan for their sponsorship and support during this project.
@@ Appendix G Additional benchmarking results
Appendix H Using other kernel functions
As described in Section 3, we used a simple rectangle function as the kernel, $K(z)$, within the pseudo-derivatives defined in Eq. 11 and Eq. 12. As shown in Fig. 17, similar results can be obtained with other common KDE kernel functions; there does not seem to be any obvious benefit to using a higher order kernel.
Figure 17: Using different kernel functions to compute the pseudo-derivatives defined in Eq. 11 and Eq. 12 has little impact on fidelity-vs-sparsity curves. These curves are for Gemma 2 9B post-layer 20 residual stream SAEs trained on 2B tokens.
Appendix I Further details on our training methodology
@@
  • As in Rajamanoharan et al. (2024), we warm up the learning rate over the first 1,000 steps (4M tokens) using a cosine schedule, starting the learning rate at 10% of its final value (i.e. starting at $7\times 10^{-6}$).
@@
Appendix J Pseudo-code for implementing and training JumpReLU SAEs
@@
### Implementation of step function with custom backward
@@
def step(x, threshold):
return (x > threshold).astype(x.dtype)
def step_fwd(x, threshold):
out = step(x, threshold)
cache = x, threshold # Saved for use in the backward pass
@@
def step_bwd(cache, output_grad):
x, threshold = cache
x_grad = zeros_like(x) # We don’t apply STE to x input
threshold_grad = (
return x_grad, threshold_grad
step.defvjp(step_fwd, step_bwd)
@@
x_grad = (x > threshold).astype(x.dtype) # We don’t apply STE to x input
@@
-(threshold / bandwidth)
* rectangle((x - threshold) / bandwidth)
* output_grad
@@
l0 = sum(step(feature_magnitudes, threshold), axis=-1)